In a class 18 children are girls and 25% of the class are boys. How many children are there altogether?
Answer by JoelSchwartz(130)

c=number of students in the class
(3/4)c=number of girls
The number of girls is equal to 18
(3/4)c=18
4(3/4)c=18*4
3c=72
c=24
(1/4)c=number of boys
.25*24=number of boys
6=number of boys